# Variable Neighborhood Searching Rerandomization

### v1

@@ B.2 Proof of Theorem 2
$Var[\widehat{\tau}]$ $=Cov\left[\beta^{T}(\overline{X}_{t}-\overline{X}_{c})\beta\right]+Var\left[\overline{e}_{t}-\overline{e}_{c}\right]$
$=\beta^{T}Cov\left[\overline{X}_{t}-\overline{X}_{c}\right]\beta+\left(1-R^{2}\right)V_{CR}$
@@
$M=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
@@
$Z=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}(W-\frac{n_{t}}{n}1_{n})=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}(\overline{X}_{t}-\overline{X}_{c}).$
@@
$Var[Z_{j}]=E[Z_{j}^{2}]=\frac{1}{p}E\left[\sum_{j=1}^{p}Z_{j}^{2}\right]=\frac{1}{p}E[M]\leq\frac{a}{p}.$
@@
Therefore, $Cov[Z]=Var[Z_{j}]I_{p}$, and
$\beta^{T}Cov[\overline{X}_{t}-\overline{X}_{c}]\beta$ $=\beta^{T}Cov\left[\sqrt{\frac{n_{t}n_{c}}{n}}S_{XX}^{1/2}Z\right]\beta$
$=\frac{n_{t}n_{c}}{n}\beta^{T}S_{XX}^{1/2}Cov[Z]S_{XX}^{1/2}\beta$
$=\frac{n_{t}n_{c}}{n}Var[Z_{j}]\beta^{T}S_{XX}\beta$
$=Var[Z_{j}]R^{2}V_{CR}$
$\leq\frac{a}{p}R^{2}V_{CR}.$
@@
$Var[\widehat{\tau}]\leq\frac{a}{p}R^{2}V_{CR}+(1-R^{2})V_{CR}.$
@@
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
@@ B.3 Proof of the Propositions
This implies that $1-W\in\mathcal{F}_{0}$ as well. Thus, $\mathcal{F}_{0}=1-\mathcal{F}_{0}$. Since CR, ARSRR, PSRR, and VNSRR treat $\mathcal{F}_{0}$ as the feasible set, they all satisfy Property 1.
StratVNSRR and ClustVNSRR conduct rerandomization with more constraints imposed by the structure of the experiments. We show that these constraints are also symmetric in nature.
We denote the feasible set of StratVNSRR and ClustVNSRR as $\mathcal{F}_{Strat}$ and $\mathcal{F}_{Clust}$, respectively.
@@
Therefore, both StratVNSRR and ClustVNSRR satisfy Property 1.
@@
For VNSRR and its variants, $(M(w^{(0)}),\ldots,M(w^{(t-1)}))$ determines whether $W^{(t)}$ is obtained through the local search step or the shaking step.
@@
And in the shaking step, $u_{t}$ is a constant which depends on the number of adjacent feasible solutions.
Therefore, VNSRR satisfies Property 3.
B.4 Theoretical Results for SeqVNSRR
In sequential experiments, we have the following Condition 4, which corresponds to the Condition 3 in the non-sequential settings.
Condition 4
@@ B.4 Theoretical Results for SeqVNSRR
We also have the corresponding version of Property 1-3, stated as follows.
Property 4
@@
Theorem 3
If $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then $E[\widehat{\tau}]=\tau$.
Theorem 4
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a_{K}}{p}\right)R^{2},$
We omit the proofs of Theorem 3 and Theorem 4, since they are almost the same as the proofs of Theorem 1 and Theorem 2, except that the chain $(W^{(0)},\ldots,W^{(T)})$ needs to be decomposed into $K$ parts corresponding to the $K$ stages. Each part can be analyzed in the same way as in simple randomized experiments.
We have verified that VNSRR satisfies Property 1-3. We note that SeqVNSRR can be viewed as a special version of VNSRR in each of its stages. As a result, SeqVNSRR satisfies the sequential version of Property 1-3.
Proposition 7
SeqVNSRR satisfies Property 5 and Property 6. Under Condition 4, SeqVNSRR satisfies Property 4.
The unbiasedness and variance reduction properties of SeqVNSRR are then immediately evident as corollaries of Theorem 3 and Theorem 4.
@@ Appendix C Randomization-based Inference
As pointed out by Zhu and Liu (2022), randomization-based confidence intervals can be constructed by inverting Fisher randomization tests. To construct a lower confidence bound for $\tau$, we consider testing $H_{0}:Y_{i}(1)-Y_{i}(0)=\theta,i=1,\ldots,n$, versus $H_{1}:Y_{i}(1)-Y_{i}(0)>\theta,i=1,\ldots,n$. Similar to testing the sharp null hypothesis with $\theta=0$, we can impute all the potential outcomes under $H_{0}$. We sample $B=1000$ acceptable assignments $\{W^{b}\}_{1\leq b\leq B}$ independently and calculate the corresponding treatment effect estimators $\{\widehat{\tau}(W^{b},\theta)\}_{1\leq b\leq B}$ using the imputed potential outcomes. The p-value is then calculated as $pv(\theta)=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b},\theta)|\geq|\widehat{\tau}(W^{obs})|\}$. Luo et al. (2021) proved that $\theta_{l}=\sup\{\theta:pv(\theta)\leq\alpha\}$ is a lower confidence bound for $\tau$ with confidence level $\alpha$. And Zhu and Liu (2022) proposed an efficient approach to solve for $\theta_{l}$, which we adopt in our numerical experiments. Similarly, we can construct an upper confidence bound $\theta_{u}$ with confidence level $\alpha$. $[\theta_{l},\theta_{u}]$ is then a confidence interval for $\tau$ with confidence level $1-2\alpha$.
@@ Appendix D Simulation Results
In Section D.1, we provide the complete results of the simulation studies in the main text. In Section D.2, we examine the randomness of assignments generated by each method. In Section D.3, we generate semi-synthetic data based on the clinical trial dataset and compare the statistical performance of each method.
@@ D.1 Statistical and Computational Performances
Table 3 and Table 4 show the complete results of our simulation studies for simple experiments and structured experiments, respectively. The results of ARSRR in the last setting of Table 3 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. The overall findings are similar to the conclusion in the main text. VNSRR substantially reduces the computational burden of rerandomization and maintains the appealing statistical properties of rerandomization.
@@ D.2 Randomness Metric
Similar to Zhu and Liu (2022), we denote $L_{n}$ as the largest eigenvalue of the covariance matrix of $12W-$ 1 and use $L_{n}$ to measure the randomness of assignments generated by each method. As pointed out by Zhu and Liu (2022), larger $L_{n}$ implies less randomness.
Table 5 and Table 6 show the results in simple experiments and structured experiments, respectively. Similar to Table 3, the results of ARSRR in the last setting of Table 5 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. CR is the most random method as expected. The three rerandomization methods perform similarly in all settings. Therefore, one can safely perform randomization-based statistical inference based on the assignments generated by VNSRR without the risk of relying on a small set of unique assignments.
D.3 Semi-synthetic Data
We mimic the set-ups in Zhu and Liu (2022) and simulate semi-synthetic data based on the clinical trial dataset. Since we only observe half of the potential outcomes in the original dataset, we impute the counterfactual potential outcomes under the sharp null. We then compare VNSRR and the existing methods in the same way as in our simulation studies. The results are shown in Table 7 and are similar to our findings in the simulation studies. VNSRR significantly accelerates rerandomization and has comparable performances in other aspects compared to ARSRR and PSRR.
Table 3: Statistical and computational performance of different methods in simple randomized experiments.
@@ D.3 Semi-synthetic Data
Bias SD Size Power Coverage Length
@@
VNSRR 5.7 181 9.8 92 90.2 608 1.3
Table 4: Statistical and computational performance of different methods in structured randomized experiments.
@@
Bias SD Size Power Coverage Length
@@
CR ARSRR PSRR VNSRR
@@ 50 $10^{-3}$ 170 231 234 233
CR ARSRR PSRR VNSRR
@@
Table 7: Statistical and computational performance of different methods when applied to semi-synthetic data.
@@
Bias SD Size Coverage Length

### v2

@@ B.2 Proof of Theorem 4
$\textnormal{Var}[\widehat{\tau}]$ $=\textnormal{Cov}\left[\beta^{T}(\overline{X}_{t}-\overline{X}_{c})\beta\right]+\textnormal{Var}\left[\overline{e}_{t}-\overline{e}_{c}\right]$
$=\beta^{T}\textnormal{Cov}\left[\overline{X}_{t}-\overline{X}_{c}\right]\beta+\left(1-R^{2}\right)\textnormal{V}_{\textnormal{CR}}$
@@
$M=\frac{n}{n_{t}n_{c}}\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}XS_{XX}^{-1}X^{T}\left(W-\frac{n_{t}}{n}1_{n}\right).$
@@
$Z=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}\left(W-\frac{n_{t}}{n}1_{n}\right)=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}(\overline{X}_{t}-\overline{X}_{c}).$
@@
$\textnormal{Var}[Z_{j}]=E[Z_{j}^{2}]=\frac{1}{p}E\left[\sum_{j=1}^{p}Z_{j}^{2}\right]=\frac{1}{p}E[M]\leq\frac{a}{p}.$
@@
Therefore, $\textnormal{Cov}[Z]=\textnormal{Var}[Z_{j}]I_{p}$, and
$\beta^{T}\textnormal{Cov}[\overline{X}_{t}-\overline{X}_{c}]\beta$ $=\beta^{T}\textnormal{Cov}\left[\sqrt{\frac{n_{t}n_{c}}{n}}S_{XX}^{1/2}Z\right]\beta$
$=\frac{n_{t}n_{c}}{n}\beta^{T}S_{XX}^{1/2}\textnormal{Cov}[Z]S_{XX}^{1/2}\beta$
$=\frac{n_{t}n_{c}}{n}\textnormal{Var}[Z_{j}]\beta^{T}S_{XX}\beta$
$=\textnormal{Var}[Z_{j}]R^{2}\textnormal{V}_{\textnormal{CR}}$
$\leq\frac{a}{p}R^{2}\textnormal{V}_{\textnormal{CR}}.$
@@
$\textnormal{Var}[\widehat{\tau}]\leq\frac{a}{p}R^{2}\textnormal{V}_{\textnormal{CR}}+(1-R^{2})\textnormal{V}_{\textnormal{CR}}.$
@@
$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
@@ B.3 Proof of the Propositions
This implies that $1-W\in\mathcal{F}_{0}$ as well. Thus, $\mathcal{F}_{0}=1-\mathcal{F}_{0}$. Since CR, ARSRR, PSRR, and BRAIN treat $\mathcal{F}_{0}$ as the feasible set, they all satisfy Property 1.
StratBRAIN and ClustBRAIN conduct rerandomization with more constraints imposed by the structure of the experiments. We show that these constraints are also symmetric in nature.
We denote the feasible set of StratBRAIN and ClustBRAIN as $\mathcal{F}_{Strat}$ and $\mathcal{F}_{Clust}$, respectively.
@@
Therefore, both StratBRAIN and ClustBRAIN satisfy Property 1.
@@
For BRAIN and its variants, $(M(w^{(0)}),\ldots,M(w^{(t-1)}))$ determines whether $W^{(t)}$ is obtained through the local search step or the perturbation step.
@@
And in the perturbation step, $u_{t}$ is a constant which depends on the number of adjacent feasible solutions.
Therefore, BRAIN satisfies Property 3.
B.4 Theoretical Results for SeqBRAIN
We consider the following equal-group-size condition for sequentially randomized experiments. We note that it has the same form as the condition for stratified randomized experiments due to the similarity between the notations in the two settings.
Condition 4.
@@ B.4 Theoretical Results for SeqBRAIN
We also have the corresponding version of Properties 1-3, stated as follows.
Property 4.
@@
Theorem 5.
If $W$ is generated by a rerandomization algorithm satisfying Properties 4-6, then $E[\widehat{\tau}]=\tau$.
Theorem 6.
Under Conditions 1 and 2, if $W$ is generated by a rerandomization algorithm satisfying Properties 4-6, then
$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a_{K}}{p}\right)R^{2},$
We omit the proofs of Theorems 5 and 6, since they are similar to the proofs of Theorems 1 and 2, except that the chain $(W^{(0)},\ldots,W^{(T)})$ needs to be decomposed into $K$ parts corresponding to the $K$ stages. Each part can be analyzed in the same way as in simple randomized experiments.
We have verified that BRAIN satisfies Properties 1-3. We note that SeqBRAIN can be viewed as a special version of BRAIN in each of its stages. As a result, SeqBRAIN satisfies the sequential version of Properties 1-3.
Proposition 4.
SeqBRAIN satisfies Properties 5 and 6. Under Condition 4, SeqBRAIN satisfies Property 4.
The unbiasedness and variance reduction properties of SeqBRAIN are then immediately evident as corollaries of Theorems 5 and 6.
@@ Appendix C Randomization-based Inference
As pointed out by [46], randomization-based confidence intervals can be constructed by inverting Fisher randomization tests. To construct a lower confidence bound for $\tau$, we consider testing $H_{0}:Y_{i}(1)-Y_{i}(0)=\theta,i=1,\ldots,n$, versus $H_{1}:Y_{i}(1)-Y_{i}(0)>\theta,i=1,\ldots,n$. Similar to testing the sharp null hypothesis with $\theta=0$, we can impute all the potential outcomes under $H_{0}$. We sample $B=1000$ acceptable assignments $\{W^{b}\}_{1\leq b\leq B}$ independently and calculate the corresponding treatment effect estimators $\{\widehat{\tau}(W^{b},\theta)\}_{1\leq b\leq B}$ using the imputed potential outcomes. The p-value is then calculated as $pv(\theta)=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b},\theta)|\geq|\widehat{\tau}(W^{obs})|\}$. [25] proved that $\theta_{l}=\sup\{\theta:pv(\theta)\leq\alpha\}$ is a lower confidence bound for $\tau$ with confidence level $\alpha$. And [46] proposed an efficient approach to solve for $\theta_{l}$, which we adopt in our numerical experiments. Similarly, we can construct an upper confidence bound $\theta_{u}$ with confidence level $\alpha$. $[\theta_{l},\theta_{u}]$ is then a confidence interval for $\tau$ with confidence level $1-2\alpha$.
@@ Appendix D Simulation Results
In D.1, we provide the complete results of the simulation studies in the main text. In D.2, we examine the randomness of assignments generated by each method. In D.3, we compare different combinations of hyperparameters. In D.4, we generate semi-synthetic data based on the clinical trial dataset and compare the statistical performance of each method.
@@ D.1 Statistical and Computational Performances
Table 3 and Table 4 show the complete results of our simulation studies for simple experiments and structured experiments, respectively. The results of ARSRR in the last setting of Table 3 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. The overall findings are similar to the conclusion in the main text. BRAIN substantially reduces the computational burden of rerandomization and maintains the appealing statistical properties of rerandomization.
@@ D.2 Randomness Metric
Similar to [46], we denote $L_{n}$ as the largest eigenvalue of the covariance matrix of $12W-$ 1 and use $L_{n}$ to measure the randomness of assignments generated by each method. As pointed out by [46], larger $L_{n}$ implies less randomness.
Table 5 and Table 6 show the results in simple experiments and structured experiments, respectively. Similar to Table 3, the results of ARSRR in the last setting of Table 5 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. CR is the most random method as expected. The three rerandomization methods perform similarly in all settings. Therefore, one can safely perform randomization-based statistical inference based on the assignments generated by BRAIN without the risk of relying on a small set of unique assignments.
D.3 Hyperparameters
There are $n$ units, hence we can swap at most $n/2$ disjoint pairs in an iteration. So, the maximum value of $L$ and $S$ is $n/2$. The minimum value of $L$ and $S$ is 1. We also consider the maximum value times 0.25, 0.5, and 0.75. Overall, there are $5\times 5=25$ combinations in total.
Figure 1 shows the computational time of BRAIN with different combinations of $L$ and $S$ under the setting where $(n,p,p_{a})=(30,15,10^{-3})$. It demonstrates that our choice of $(L,S)$ is among the top tier.
D.4 Semi-synthetic Data
We mimic the set-ups in [46] and simulate semi-synthetic data based on the clinical trial dataset. Since we only observe half of the potential outcomes in the original dataset, we impute the counterfactual potential outcomes under the sharp null. We then compare BRAIN and the existing methods in the same way as in our simulation studies. The results are shown in Table 7 and are similar to our findings in the simulation studies. BRAIN significantly accelerates rerandomization and has comparable performances in other aspects compared to ARSRR and PSRR.
Table 3: Statistical and computational performance of different methods in simple randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
@@ D.4 Semi-synthetic Data
Bias SD Size Power CP Len
@@
BRAIN 5.7 181 9.8 92 90.2 608 1.3
Table 4: Statistical and computational performance of different methods in structured randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
@@
Bias SD Size Power CP Len
@@
CR ARSRR PSRR BRAIN
@@ 50 $10^{-3}$ 170 231 234 233
CR ARSRR PSRR BRAIN
@@
Figure 1: Comparison of different hyperparameters.
Table 7: Statistical and computational performance of different methods when applied to semi-synthetic data. SD: standard deviation. CP: coverage probability. Len: length.
@@
Bias SD Size CP Len